OpenStudy (waheguru):

when multiplying numbers is it true that the largest number kepps its sigh whether it being a negative or a positive PLz ANSWER FAST NEED HELP

OpenStudy (anonymous):

sure

OpenStudy (waheguru):

yea or not not sure

OpenStudy (anonymous):

It depends on the sign of the other number.

OpenStudy (waheguru):

plzzzzzzzzzzzzzzzzzzzzzzz answer guys [

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so if you have -1 times 1 then it is -1 1times 1 is 1 and -1times -1 is 1

OpenStudy (anonymous):

Counter example: +100000 * -3 = - 300000 It seems that the largest number does not keep its sign

OpenStudy (waheguru):

oh

OpenStudy (anonymous):

do you get it?

OpenStudy (waheguru):

so then how do i do long multiplaction with a negative and a psitive number

OpenStudy (anonymous):

the negetive wins

OpenStudy (anonymous):

- times + then it is - 1times + is + and - times - is +

OpenStudy (lalaly):

only when adding u take the sign of the largest number like -4+2 = -2

OpenStudy (waheguru):

so if there is i negative the whole answer becomes negative

OpenStudy (anonymous):

- times + then it is - + times + is + and - times - is +

OpenStudy (anonymous):



OpenStudy (anonymous):

there are some examples

OpenStudy (waheguru):

thx izzy25

OpenStudy (anonymous):

NP, keep it,write it down somewhere it helps

OpenStudy (anonymous):

(-)*(-)=(+) (-)*(+)=(-) (+)*(-)=(-) (+)*(+)=(+)